# Attention Mechanisms Don’t Learn Additive Models: Rethinking Feature Importance for Transformers

### v1

@@ 5.4 Relating SLALOM scores to linear attributions
Importantly, SLALOM scores can be readily converted to locally linear interpretability scores for applications where necessary. For this purpose, a differentiable model for soft removals is required. We consider the following weighted model:
@@
where $\lambda_{i}=1$ if a token is present and $\lambda_{i}=0$ if it is absent. We observe that setting $\lambda_{i}=0$ has the desired effect of making the output of the soft-removal model equivalent to that of the standard SLALOM on a sequence without this token. Taking the gradients at ${\bm{\lambda}}=\mathbf{1}$ we obtain $\left.\frac{\partial F}{\partial\lambda_{i}}\right\rvert_{{\bm{\lambda}}=\mathbf{1}}\propto v(t_{i})\exp(s(t_{i}))$, which can be used to rank tokens according the linearized attributions. We defer the derivation to Section B.7 and refer to these scores as linearized SLALOM scores.
@@ 6 Experimental Evaluation
We run a series of experiments show the mismatch between surrogate model explanations and the transformers. Specifically, we verify that (1) real transformers fail to learn additive models, (2) SLALOM better captures transformer output, (3) SLALOM models can be reidentified from fitted models with tolerable error, and (4) SLALOM is more versatile than other methods as it aligns will with linear attribution scores and human attention, and performs better in faithfulness metrics. For experiments (1)-(3), we require knowledge of the ground truth, and therefore use synthetic datasets. To demonstrate the practical strengths of our method, all the experiments for (4) are conducted on real-world datasets, in comparison with state-of-the-art XAI techniques.
@@ 6.1 Experimental Setup
LM architectures. We study three representative transformer language model architectures in our experiments. In sequence classification, mid-sized transformer LMs are most common  [24], which is reflected by our experimental setup. To represent the family of encoder-only models, we deploy BERT [15] and DistilBERT [43]. We further experiment with GPT-2 [39], which is a decoder-only model. We use the transformers framework to run our experiments. While not the main scope of this work, we show that SLALOM can be applied to LLMs with up to 7B parameters and non-transformer models such as Mamba [21] with plausible results due to its general expressivity in Section F.8.
Datasets. We use two real-world datasets for sentiment classification. Specifically, we study the IMDB dataset, consisting of movie reviews [32] and Yelp-HAT, a dataset of Yelp reviews, for which human annotators have provided annotations on which tokens are relevant for the classification outcome [45]. We provide additional details on training and datasets in Appendix E.
@@ 6.2 Evaluation with Known Ground Truth
Transformers fail to capture linear relationships. We start our experimental section by experimentally verifying the claims made in Proposition 4.1 and Corollary 4.2. To ascertain that the underlying relation captured by the models is additive, we resort to a synthetic dataset. The dataset is created as follows: First, we sample different sequence lengths from a binomial distribution with a mean of $15$. Second, we sample words independently from a vocabulary of size $10$. This vocabulary was chosen to include positive words, negative words, and neutral words, with manually assigned weights $w\in\{-1.5,-1,0,1,1.5\}$, that can be used to compute a linear log-odds model. We evaluate this model and finally sample the sequence label accordingly, thereby ensuring a linear relation between input sequences and log-odds. We train transformer models on this dataset and evaluate them on sequences containing the same word (“perfect”) multiple times. Our results in Figure 3 show that the models fail to capture the relationship regardless of the model or number of layers used. In Appendix A, we show how this undermines the recovery property with Shapley value explanations.
Fitting SLALOM as a surrogate to transformer models. Having demonstrated the mismatch between additive functions and transformers, we turn to SLALOM as a more suitable surrogate model. As shown in Proposition 5.1, transformers can easily fit SLALOMs, which is why we hypothesize that they should model the output of such a model well in practice. We observe that this is indeed the case in Figure 4(a, b) even if the GAM has considerably more parameters than SLALOM.
Verifying Recovery. We run an experiment to study whether, unlike linear models, SLALOM can be fitted and recovered by transformers. To test this, we sample a dataset that exactly follows the distribution given by SLALOM. We then train transformer models on this dataset. The results in Figure 4(c, d) show that the surrogate model fitted on transformer outputs as a post-hoc explanation recovers the correct log-odds mandated by SLALOM (c) and that there is a good correspondence between the true model’s and the recovered model’s parameters (d).
@@
Figure 5: Explaining a real review with SLALOM (qualitative results). SLALOM assigned two scores to each token and can be used to compute attributions. See Figure 11 for fully annotated plots.
@@ 6.3 Examining Real-World Predictions from Different Angles
We increase difficulty and deploy SLALOM to explain predictions on real-world datasets. As there is no ground truth for these datasets, it is challenging to evaluate the quality of the explanations [41]. However, we provide comparisons from several angles: We compare to linear scores obtained when fitting a Naïve-Bayes Bag-of-Words model, scores on removal and insertion benchmarks [53, 16], the human attention scores available on the Yelp-HAT dataset [45], and provide qualitative results.
Explaining Sentiment Classification. We show qualitative results for explaining a movie review in Figure 5. The figure shows that both negative and positive words are assigned high importance scores but have value scores of different signs. Furthermore, we see that some words (“the”) have positive value scores, but a very low importance. This means that they lead to positive scores on their own but are easily overruled by other words. We compare the SLALOM scores obtained on $50$ random test samples to a linear Naïve-Bayes model (obtained though counting class-wise word frequencies) as a surrogate ground truth in Table 2(a) through the Spearman rank correlation. We observe good agreement with the value scores $v$ and the linearized SLALOM scores (“lin”, see Section 5.4).
Evaluating Faithfulness Metrics. Faithfulness metrics quantify the alignment of the explanations with actual model behavior and can be used to compare different explanation techniques. We run the classical removal benchmarks with SLALOM compared to baselines such as LIME [40], SHAP [31], Gradients [48], and Integrated Gradients (IG, [51]). For the deletion benchmark, the tokens with the highest attributions are successively deleted from the sample which should result in a low score for the target class. We iteratively delete more tokens and compute the “Area Over the Perturbation Curve” (AOPC, see [16]), which should be high for deletion. For surrogate techniques (LIME, SHAP, SLALOM) we use 5000 samples each. Our results in Tab. 2(c) highlight that linearized SLALOM scores outperform LIME and SHAP. Insertion results with similar findings are deferred to Table 9.
@@
LM values $v$ importances $s$ lin. LIME SHAP
D-BERT 0.68 $\pm$ 0.03 0.73 $\pm$ 0.02 0.71 $\pm$ 0.03 0.67 $\pm$ 0.03 0.69 $\pm$ 0.03
BERT 0.57 $\pm$ 0.04 0.72 $\pm$ 0.03 0.64 $\pm$ 0.03 0.63 $\pm$ 0.03 0.68 $\pm$ 0.03
GPT-2 0.67 $\pm$ 0.02 0.67 $\pm$ 0.02 0.67 $\pm$ 0.02 0.64 $\pm$ 0.03 0.60 $\pm$ 0.03
@@
(b) Measuring average AU-ROC between SLALOM explanations and human token attention. The importance scores are clearly and most strongly predictive of human attention.
LM values importances $s$ lin. LIME SHAP IG Grad
DistilBERT 0.913 $\pm$ 0.053 0.249 $\pm$ 0.215 0.915 $\pm$ 0.053 0.892 $\pm$ 0.068 0.876 $\pm$ 0.071 0.360 $\pm$ 0.346 0.387 $\pm$ 0.344
BERT 0.943 $\pm$ 0.048 0.262 $\pm$ 0.252 0.944 $\pm$ 0.048 0.939 $\pm$ 0.056 0.929 $\pm$ 0.070 0.093 $\pm$ 0.130 0.093 $\pm$ 0.129
GPT-2 0.907 $\pm$ 0.055 0.294 $\pm$ 0.154 0.908 $\pm$ 0.055 0.905 $\pm$ 0.057 0.895 $\pm$ 0.063 0.378 $\pm$ 0.203 0.436 $\pm$ 0.204
@@
(c) Area Over Perturbation Curve for deletion. Linearized scores performs best in the faithfulness metric.
@@
Table 2: Evaluation of SLALOM scores (“values”, “importance”, “lin.”) with std. errors across explanation quality measures highlights that SLALOM’s different scores serve different purposes.
@@
Predicting Human Attention. To incorporate a user perspective we predict human attention from SLALOM scores. We compute AU-ROC for predicting annotated human attention as suggested in Sen et al. [45] in Table 2(b). We use absolute values of all signed explanations as human attention is unsigned as well. In contrast to the previous experiments, where linearized scores and value scores were more effective, we observe that the importance scores are better at predicting where the human attention is placed, and are more effective than LIME and SHAP. We omit Grad and IG, which perform almost on random level for this task. In summary, these findings highlight the fact that different tasks require different attribution scores and that a single score is insufficient to explain transformer models, as suggested by our theory. SLALOM offers flexibility through its $2$-dimensional representation and provides state-of-the-art and efficient attribution explanations on real-world tasks.
@@ 7 Discussion and Conclusion
In this work, we established that transformer networks are inherently incapable of aligning with linear or additive models for feature attribution. This may explain similar incapacities observed in time-series forecasting [59], where they seem incapable of representing certain relations. To address this shortcoming, we have introduced the Softmax-Linked Additive Log-Odds Model (SLALOM), a surrogate model for explaining the influence of features on transformers and other complex LMs through a two-dimensional representation. To complement this theoretical foundation, future work will include the evaluation of SLALOM from a user-centric perspective, for instance, on human-centered evaluation frameworks [13]. From a broader perspective, we hope that this research paves the way for advancing the interpretability of widely adopted transformer models.

### v2

@@ 5.4 Relating SLALOM scores to linear attributions
Importantly, SLALOM scores can be readily converted to locally linear interpretability scores where necessary. For this purpose, a differentiable model for soft removals is required. We consider the weighted model:
@@
where $\lambda_{i}=1$ if a token is present and $\lambda_{i}=0$ if it is absent. We observe that setting $\lambda_{i}=0$ has the desired effect of making the output of the soft-removal model equivalent to that of the standard SLALOM on a sequence without this token. Taking the gradients at ${\bm{\lambda}}=\mathbf{1}$ we obtain $\left.\frac{\partial F}{\partial\lambda_{i}}\right\rvert_{{\bm{\lambda}}=\mathbf{1}}\propto v(t_{i})\exp(s(t_{i}))$, which can be used to rank tokens according to the linearized attributions. We defer the derivation to Section B.7 and refer to these scores as linearized SLALOM scores. As SLALOM is just another multi-variate function, it also possible to compute SHAP values for it. We discuss this relation in Section B.7 as well.
@@ 6 Experimental Evaluation
We run a series of experiments to show the mismatch between surrogate model explanations and the transformers. Specifically, we verify that (1) real transformers fail to learn additive models, (2) SLALOM better captures transformer output, (3) SLALOM models can be recovered from fitted models with tolerable error, (4) SLALOM scores are versatile and align well with linear attribution scores and human attention, and (5) that SLALOM performs well in faithfulness metrics and has substantially higher fidelity than competing techniques. For experiments (1)-(3), we require knowledge of the ground truth and use synthetic datasets. To demonstrate the practical strengths of our method, all the experiments for (4) and (5) are conducted on real-world datasets and in comparison with state-of-the-art XAI techniques.
@@ 6.1 Experimental setup
LM architectures. We study three representative transformer language model architectures in our experiments. In sequence classification, mid-sized transformer LMs are most popular on platforms such as the Huggingface hub  (Huggingface, 2023) often based on the BERT-architecture (Devlin et al., 2018), which is reflected in our experimental setup. To represent the family of encoder-only models, we deploy BERT (Devlin et al., 2018) and DistilBERT (Sanh et al., 2019). We further experiment with GPT-2 (Radford et al., 2019), which is a decoder-only model. We use the transformers framework to run our experiments. While not the main scope of this work concerned with sequence classification, we will also show that due to its model-agnostic nature, SLALOM can be applied to LLMs with up to 7B parameters and non-transformer models such as Mamba (Gu & Dao, 2023) with plausible results. We provide a proof-of-concept to show that SLALOM can be applied to black-box models such as GPT-4o in Section F.7.
Datasets. We use synthetic datasets and two real-world datasets for sentiment classification. Specifically, we study the IMDB dataset, consisting of movie reviews (Maas et al., 2011) and Yelp-HAT, a dataset of restaurant reviews, for which human annotators have provided annotations on which tokens are relevant for the classification outcome (Sen et al., 2020). We provide additional details on hyperparameters, training, and datasets in Appendix E.
@@ 6.2 Evaluation with known ground truth
Transformers fail to capture linear relationships. We empirically verify the claims made in Proposition 4.1 and Corollary 4.3. To ascertain that the underlying relation captured by the models is additive, we resort to a synthetic dataset with a linear relation. The dataset is created as follows: First, we sample different sequence lengths from a binomial distribution with a mean of $15$. Second, we sample words independently from a vocabulary of size $10$. This vocabulary was chosen to include positive words, negative words, and neutral words, with manually assigned weights $w\in\{-1.5,-1,0,1,1.5\}$, that can be used to compute a linear log odds model. We evaluate this model and finally sample the sequence label accordingly, thereby ensuring a linear relation between input sequences and log odds. We train transformer models on this dataset and evaluate them on sequences containing the same word (“perfect”) multiple times. Our results in Figure 3 show that the models fail to capture the relationship regardless of the model or number of layers used. In Appendix A, we show how this undermines the recovery property with Shapley value explanations.
Fitting SLALOM as a surrogate to transformer models. Having demonstrated the mismatch between additive functions and transformers, we turn to SLALOM as a more suitable surrogate model. As shown in Proposition 5.1, transformers can easily fit SLALOMs, which is why we hypothesize that they should model the output of such a model well in practice. We fit the different surrogate models on a dataset of input sequences and real transformer outputs from our linear synthetic dataset and observe that linear models and additive models fail to capture the relationship learned by the transformer as shown in Figure 4(a, b). On the contrary, SALO manages to model the relationship well, even if it has considerably less trainable parameters than the additive model (GAM).
Verifying recovery. We run an experiment to study whether, unlike linear models, SLALOM can be fitted and recovered by transformers. To test this, we sample a second synthetic dataset that exactly follows the relation given by SLALOM. We then train transformer models on this dataset. The results in Figure 4(c, d) show that the surrogate model fitted on transformer outputs as a post-hoc explanation recovers the correct log odds mandated by SLALOM (c) and that there is a good correspondence between the true model’s parameters and the recovered model’s parameters (d).
@@
Figure 5: Explaining a real review with SLALOM (qualitative results). SLALOM assigns two scores to each token (a,b) and can be used to compute attributions via its linearization (c). We observe that the impactful words have high importances and the value scores indicate the sign of their contribution (positive or negative words). See Figure 12 (Appendix) for fully annotated plots.
@@ 6.3 Examining real-world predictions from different angles
We increase the difficulty and deploy SLALOM (fitted using SLALOM-eff) to explain predictions on real-world datasets. As there is no ground truth for these datasets, it is challenging to evaluate the quality of the explanations (Rong et al., 2022). To better understand SLALOM explanations, we study them from several angles: We compare to linear scores obtained when fitting a Naïve-Bayes Bag-of-Words (BoW) model, scores on removal and insertion benchmarks (Tomsett et al., 2020; DeYoung et al., 2020), the human attention scores available on the Yelp-HAT dataset (Sen et al., 2020), and provide qualitative results.
Explaining Sentiment Classification. We show qualitative results for explaining a movie review in Figure 5. The figure shows that both negative and positive words are assigned high importance scores but have value scores of different signs. Furthermore, we see that some words (“the”) have positive value scores, but a very low importance. This means that they lead to positive scores on their own but are easily overruled by other words. We compare the SLALOM scores obtained on $100$ random test samples to a linear Naïve-Bayes model (obtained though counting class-wise word frequencies) as a surrogate ground truth in Table 2(a) through the Spearman rank correlation. We observe good agreement with the value scores $v$ and the combined linearized SLALOM scores (“lin”, see Section 5.4).
@@
LM values $v$ importances $s$ lin.
Bert 0.786 $\pm$ 0.01 0.807 $\pm$ 0.01 0.801 $\pm$ 0.01
DistilBERT 0.688 $\pm$ 0.01 0.681 $\pm$ 0.01 0.686 $\pm$ 0.01
GPT-2 0.674 $\pm$ 0.01 0.685 $\pm$ 0.01 0.683 $\pm$ 0.01
average 0.716 0.724 0.724
BLOOM-7.1B 0.739 $\pm$ 0.02 0.712 $\pm$ 0.03 0.740 $\pm$ 0.02
Mamba-2.8B 0.615 $\pm$ 0.03 0.437 $\pm$ 0.03 0.535 $\pm$ 0.03
@@
(b) Measuring average AU-ROC between SLALOM explanations and human token attention. The importance scores and most strongly predictive of human attention for the classical models. Applying SLALOM to larger models underlines its scalability, but it remains less reliable for non-transformer models like Mamba.
SLALOM-fidel SLALOM-eff
LM $v$-scores lin. $v$-scores lin. LIME SHAP IG Grad LRP
BERT 0.025 $\pm$ 0.002 0.023 $\pm$ 0.001 0.031 $\pm$ 0.002 0.031 $\pm$ 0.002 0.024 $\pm$ 0.002 0.026 $\pm$ 0.003 0.557 $\pm$ 0.034 0.611 $\pm$ 0.033 0.030 $\pm$ 0.006
DistilBERT 0.028 $\pm$ 0.003 0.024 $\pm$ 0.002 0.027 $\pm$ 0.002 0.027 $\pm$ 0.002 0.027 $\pm$ 0.003 0.029 $\pm$ 0.003 0.495 $\pm$ 0.027 0.508 $\pm$ 0.028 0.023 $\pm$ 0.002
GPT-2 0.052 $\pm$ 0.008 0.050 $\pm$ 0.008 0.089 $\pm$ 0.008 0.089 $\pm$ 0.008 0.230 $\pm$ 0.017 0.042 $\pm$ 0.005 0.454 $\pm$ 0.022 0.493 $\pm$ 0.023 0.069 $\pm$ 0.010
average 0.035 $\pm$ 0.004 0.032 $\pm$ 0.004 0.049 $\pm$ 0.004 0.049 $\pm$ 0.004 0.094 $\pm$ 0.007 0.032 $\pm$ 0.003 0.502 $\pm$ 0.028 0.537 $\pm$ 0.028 0.041 $\pm$ 0.006
@@
(c) Area Over Perturbation Curve for deletion. Linearized scores and SHAP performs best in the XAI metric.
@@
Table 2: Evaluation of SLALOM scores (“values”, “importance”, “lin.”) with std. errors across explanation quality measures highlights that SLALOM’s different scores serve different purposes. IMDB dataset shown.
@@
Predicting Human Attention. To study alignment with a user perspective, we predict human attention from SLALOM scores. We compute AU-ROC for predicting annotated human attention as suggested in Sen et al. (2020) in Table 2(b). We use absolute values of all signed explanations as human attention is unsigned as well. In contrast to the previous experiments, where value scores were more effective than importances, we observe that the importance scores are often best at predicting where the human attention is placed. In summary, these findings highlight that the two scores serve different purposes and cover different dimensions of interpretability. SLALOM offers higher flexibility through its $2$-dimensional representation.
We also use this opportunity to study the applicability of SLALOM to larger and non-transformer models. To this end, we train Mamba and BLOOM models (Le Scao et al., 2023) with a classification head on the Yelp dataset. Our results in the lower part of Table 2(b) prove that SLALOM can be well applied to larger models with sizes of up to 7B. For BLOOM the results are promising and reach the same level as for the classical transformer models. SLALOM can also be applied to non-transformer models like Mamba (Gu & Dao, 2023) due to its model-agnostic nature. Due to its general expessivity, the explanations are capable of predicting human attention (at least the value scores), but we observe a slight reduction in quality. We thus see SLALOM’s main scope with models that follow the classical transformer paradigm. We show that SLALOM’s results for BoW correlations and human attention prediction are in the same range and often outperform competing XAI techniques in Section F.3.2, but defer a comparative analysis to the next sections.
(a) BERT
(b) DistilBERT
(c) GPT-2
Figure 6: Assessing Fidelity. We plot the MSE for predicting model outputs under token removal and find that SLALOM’s predictions have up to 70% less error than the closest competitor when up to 10 random tokens from a sentence are removed (log-$y$ plots). We interpret LRP scores as a linear model.
Assessing Fidelity. Having established the roles of its components, we verify that SLALOM can produce explanations that have substantially higher fidelity than competing surrogate or non-surrogate explanation techniques. To assess this, we remove up to 10 tokens from the input sequences and use the explanations to predict the change in model output using the surrogate model (SLALOM or linear). We compare the two SLALOM versions to baselines such as LIME (Ribeiro et al., 2016), Kernel-SHAP (Lundberg & Lee, 2017), Gradients (Simonyan et al., 2013), and Integrated Gradients (IG, Sundararajan et al., 2017) and layer-wise relevance propagation (LRP) for transformers (Achtibat et al., 2024), a non-surrogate technique. We report the Mean-Squared-Error (MSE) between the predicted change and the observed change when running the model on the modified inputs in Figure 6. We observe that SLALOM-fidel offers substantially higher fidelity with a reduction of 70% in MSE for predicting the output changes over the second-best method (LRP). Other surrogate approaches and LRP remain cluttered together, potentially highlighting the frontier of maximum fidelity possible with a linear surrogate model.
Approach Avg. Time (s)
Grad 0.01 $\pm$ 0.00
IG 0.02 $\pm$ 0.00
LRP 0.02 $\pm$ 0.00
SLALOM-eff 2.03 $\pm$ 0.01
SLALOM-fidel 3.77 $\pm$ 0.24
LIME 3.93 $\pm$ 0.19
SHAP 11.56 $\pm$ 0.03
Table 3: Runtime comparison using 5000 samples to estimate surrogate models.
Evaluating XAI Metrics. There are several other metrics to quantify the quality of explanations and to compare different explanation techniques. As a sanity check and to show that SLALOM explanations do not lag behind other techniques in established metrics, we run the classical insertion/removal benchmarks (Tomsett et al., 2020). For the insertion benchmark, we successively add the tokens with the highest attributions to the sample, which should result in a high score for the target class. We iteratively insert more tokens and compute the “Area Over the Perturbation Curve” (AOPC, see DeYoung et al. (2020)), which should be low for insertion. This metric quantifies the alignment of explanations and model behavior but only considers the feature ranking and not the assigned score. For surrogate techniques (LIME, SHAP, SLALOM) we use 5000 samples each. Our results in Tab. 2(c) highlight that linearized SLALOM scores outperform LIME and LRP and perform on par with SHAP. Removal results for IMDB and results on YELP with similar findings are deferred to Table 12 (Appendix). In conclusion, this shows that on top of SLALOM’s desirable properties, it is on par with other techiques in common evaluation metrics.
Computational Costs. Finally, we take a look at the computational costs of the methods, which are mainly determined by sampling the dataset to fit the surrogate model. We provide the runtimes to explain a sample on our hardware when using 5000 samples to estimate surrogates in Table 3 with more results in Section F.6. We observe that SHAP incurs the highest computational burden. Among surrogate model explanations SLALOM-eff is the most efficient, being about 5$\times$ more efficient that SHAP and 2$\times$ more efficient than LIME. Nevertheless, non-surrogate techniques are far more efficient as they require only one or few (IG steps) forward or backward passes, but suffer from other disadvantages (e.g., implementation effort, no explicit way to predict model behavior). Overall, our results highlight that the two SLALOM fitting routines can produce explanations of comparable utility to other surrogate models at a fraction of the costs, or produce explanations with higher fidelity at similar costs due to structurally better alignment between surrogate and predictive models.
@@ 7 Discussion and Conclusion
In this work, we established that transformer networks are inherently incapable of representing linear or additive models commonly used for feature attribution. We prove that the function spaces learnable by transformers and linear models are disjoint when ruling out trivial cases. This may explain similar incapacities observed in time-series forecasting (Zeng et al., 2023), where they seem incapable of representing certain relations. To address this shortcoming, we have introduced the Softmax-Linked Additive Log Odds Model (SLALOM), a surrogate model for explaining the influence of features on transformers and other complex LMs through a two-dimensional representation.
Our work still has certain limitations that could be addressed in future work. SLALOM is specifically designed to explain the behavior of transformer models and therefore aligned with the classes of functions commonly represented by transformers. However, it would not be a suitable choice to explain models capturing a linear relationship. While SLALOM is generally applicable to any token-based LMs, we recommend using SLALOM only when the model is known to have attention-like non-linearities. Our results indicate that performance for these models is highest. We also note that SLALOM operates at the token level by assigning each individual token importance and value scores. Contextual or higher-order interpretability considering the meaning and impact of phrases, clauses, or sentences is not covered by SLALOM. To complement this theoretical foundation, future work will include further evaluation of SLALOM from a user-centric perspective, for instance, on human-centered evaluation frameworks (Colin et al., 2022). From a broader perspective, we hope that this research paves the way for advancing the interpretability and theoretical understanding of widely adopted transformer models.
Broader Impact Statement
This paper presents theoretical work on better understanding feature attributions in the transformer framework. We advise using caution when using our XAI technique or other model explanation as all explanations present only a simplified view of the complex ML model. Our method works best with models of the transformer architecture. Besides that, we do not see any immediate impact which we feel must be specifically highlighted here.